OpenStudy (anonymous):

Prove using trig funtions: tan^2 cos^2 + cos^2 = 1 Please show steps if possible :D

OpenStudy (solomonzelman):

tan^2x cos^2x + cos^2x = 1 not just tan^2

OpenStudy (solomonzelman):

\[\tan^2x ~~\cos^2x + \cos^2x = 1 \]\[(\sin^2x/\cos^2x)~~ \cos^2x + \cos^2x = 1\]\[\sin^2x + \cos^2x = 1\]makes sense ?

OpenStudy (anonymous):

Yes! Thanks a lot. I appreciate it! will give medal

OpenStudy (solomonzelman):

If you want to give it to me, I don';t mind of course :)

OpenStudy (solomonzelman):

You welcome !